OpenStudy (anonymous):

What is the equation of a line that passes through points (0, 4) and (-4, -8)? y = 2x + 4 y = 2x - 4 y = 3x + 4 y = 3x - 4

ganeshie8 (ganeshie8):

start by finding the slope

OpenStudy (astrophysics):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (anonymous):

a

OpenStudy (astrophysics):

That was fast

ganeshie8 (ganeshie8):